# GALD-SE: Guided Anisotropic Lightweight Diffusion for Efficient Speech Enhancement

### v1

@@ III-C Results
We initially compare our method with others using the DNS corpus, as shown in Table I, where ‘D’ denotes discriminative methods and ‘G’ represents generative approaches. Our results demonstrate that the proposed method achieves performance comparable to state-of-the-art techniques across multiple evaluation metrics. Although our method shows a slight performance gap in ESTOI and SI-SNR metrics compared to the best-performing BBED approach, it achieves this with significantly fewer parameters (4.5M vs. 65M) and sampling steps (10 vs. 60). We attribute this significant computational efficiency to our novel implementation of anisotropic noise guidance, which preserves clean speech clues and streamlines the diffusion process. It is worth noting that PESQ may lack reliability in low-SNR scenarios, where ESTOI emerges as a more suitable metric. Our results demonstrate that performance improvements are substantially more pronounced in low-SNR conditions compared to high-SNR situations (an ESTOI improvement of 0.25 in the -15 to -5 dB range versus 0.14 in the 5 to 15 dB range). This underscores the efficacy of our method in challenging acoustic environments.
@@
Methods Type PESQ ESTOI CSIG CBAK COVL
Mixture - 1.97 0.79 3.35 2.44 2.63
NCSN++ [24] D 2.73 0.86 3.45 3.39 3.17
UNet [30] D 2.87 0.87 4.20 3.48 3.60
DEMUCS [37] D 3.07 - 4.31 3.40 3.63
CDiffuse [11] G 2.52 - 3.72 2.91 3.10
SGMSE+ [9] G 2.93 0.87 4.12 3.37 3.51
StoRM [12] D+G 2.93 0.88 - - -
Thunder [23] D+G 2.97 0.87 - - -
VPIDM [13] G 3.16 0.87 4.23 3.53 3.70
GALD-SE (Ours) D+G 3.19 0.88 4.25 3.59 3.73
   –w/o guidance G 2.93 0.85 4.10 3.15 3.48
   –w/o adding noise G 2.79 0.83 3.97 2.99 3.35
Subsequently, we evaluate our method against other state-of-the-art approaches on the VBD dataset. The results are presented in Table II. The NCSN++ approach adapts the diffusion model architecture to function as a discriminative backbone, whereas the UNet approach exclusively employs the first-stage coarse magnitude estimation, leaving the phase unaltered. Our results demonstrate that the combination of diffusion with anisotropic guidance outperforms approaches using either the coarse magnitude frontend or diffusion as a discriminative backbone alone. Moreover, when compared to other state-of-the-art methods, our proposed approach achieves superior performance across most of the metrics. We posit that this advantage stems from our novel strategy of deeply integrating clean speech clues into the diffusion process through guidance, which enables a more nuanced and effective speech enhancement process.
Finally, we conducted an ablation study to demonstrate the effectiveness of our anisotropic guidance method, as shown in the last three rows of Table II. We compared our proposed method against two variants: One without anisotropic guidance and another without adding Gaussian noise entirely. The latter case, where Gaussian noise is absent throughout the process, corresponds to the cold diffusion formulation [26]. Results indicate that incorporating an appropriate amount of Gaussian noise during the process, rather than directly corrupting clean speech with original noise, yields better results. Moreover, applying guidance to the Gaussian noise during diffusion significantly enhances performance across all evaluated metrics, thus validating the effectiveness of our proposed method.

### v5

@@ III-C Results
We initially compare our method with others using the simulated DNS corpus, as presented in Table I, where ‘D’ and ‘G’ denote discriminative and generative approaches, respectively. To ensure a comprehensive evaluation, we select representative methods from both categories, encompassing a range of established and state-of-the-art techniques. For the discriminative approach, we choose widely-adopted advanced methods, including DEMUCS [38], CMGAN [39], and DeepFilterNet [40]. Regarding the generative approach, we focus on recent diffusion-based methods, such as SGMSE+ [9], StoRM [12], VPIDM [13] and BBED [16]. Our results demonstrate that the proposed method achieves performance comparable to state-of-the-art techniques across multiple evaluation metrics. Although our method shows a slight performance gap in ESTOI and SI-SNR metrics compared to the best-performing BBED approach, it achieves this with significantly fewer parameters (4.5M vs. 65M) and sampling steps (6 vs. 60). We attribute this significant computational efficiency to our novel implementation of anisotropic noise guidance, which preserves clean speech clues and streamlines the diffusion process. It is worth noting that PESQ may lack reliability in low-SNR scenarios, where ESTOI emerges as a more suitable metric. Our results demonstrate that our method achieves an ESTOI rank of No. 2 among all methods, despite having significantly fewer parameters than the top-ranked BBED method. This underscores the efficacy of our approach in challenging acoustic environments.
@@
Methods Type Para. Steps PESQ ESTOI CSIG CBAK COVL
Mixture - - - 1.97 0.79 3.35 2.44 2.63
NCSN++ [25] D 3.6M - 2.73 0.86 3.45 3.39 3.17
UNet-Small [30] D 0.9M - 2.85 0.86 3.97 3.43 3.52
DEMUCS [38] D 33.5M - 3.07 - 4.31 3.40 3.63
CDiffuSE [11] G 10.1M 6 2.52 - 3.72 2.91 3.10
SGMSE+ [9] G 65M 30*2 2.93 0.87 4.12 3.37 3.51
StoRM [12] D+G 55.6M 15*2 2.93 0.88 - - -
Thunder [23] D+G 27.8M 30*2 2.97 0.87 - - -
VPIDM [13] G 65M 25*2 3.16 0.87 4.23 3.53 3.70
GALD-SE (Ours) D+G 4.5M 6 3.19 0.88 4.23 3.61 3.72
–reduced CMEN D+G 3.9M 6 3.15 0.88 4.21 3.57 3.72
–with isotropic noise D+G 4.5M 6 2.93 0.87 4.10 3.26 3.50
–noise-free D+G 4.5M 6 2.80 0.83 4.02 3.11 3.38
Subsequently, we compare our method with others on the VBD dataset. The results are presented in Table II. We include the recently proposed Thunder [23] in this evaluation, as it was originally assessed on the VBD dataset, enabling direct performance comparison. The NCSN++ approach adapts the diffusion model architecture to function as a discriminative backbone, whereas the UNet approach exclusively employs the first-stage coarse mask estimation, leaving the phase unaltered. Our results demonstrate that the combination of diffusion with anisotropic guidance outperforms approaches using either the coarse mask frontend or diffusion as a discriminative backbone alone. Moreover, when compared to other state-of-the-art methods, our proposed approach achieves superior performance across most of the metrics. We posit that this advantage arises from our novel strategy of deeply integrating clean speech clues into the diffusion process through guidance, which enables a more nuanced and effective speech enhancement process.
Finally, we conduct an ablation study to demonstrate the effectiveness of our anisotropic guidance method, as shown in the last four rows of Table II. We evaluate three variants: the first adopting a reduced CMEN architecture with two encoder-decoder layers, the second incorporating isotropic noise during diffusion, and the third implementing a noise-free process. The last case corresponds to the cold diffusion formulation [27]. For fair comparison, all variants incorporate the discriminative guidance $\boldsymbol{\sigma}$ as input to the diffusion network. Results from the first variant show only marginal performance degradation with reduced CMEN parameters, demonstrating our framework’s robustness. The isotropic noise and noise-free variants, however, exhibit evident performance deterioration, validating the necessity of guided anisotropic noise in our method.